OpenStudy (anonymous):

Help me please with this integral (x^2)(e^(-x^4/28))dx. I know I have to do it for parts but I get stuck when I have to integrate e :(

OpenStudy (anonymous):

This doesn't look like it has a nice elementary solution... Are you sure it's \(x^4\) in the exponent?

OpenStudy (anonymous):

Yes. I'm solving \[\frac{ dy }{ dx } - \frac{ x^3 }{ 7 }(y)=4x^2\] by integral factor and in my result I have to integrate \[\int\limits 4x^2 e^{-x^4/28} dx\]

OpenStudy (anonymous):

Ah, if this is for solving a differential equation, you can leave your solution in this non-closed form.

OpenStudy (anonymous):

Thank you, I think I'll do that!

OpenStudy (anonymous):

yw